please help.
The length of a rectangle in 9in. more than twice its width. If the peremiter of the rectangle is 48in.,find the width of the rectangle.

Answer by neta(14)

the eqn for the problem is
let the width be x, the perimeter is give as 48
perimeter=2(l+b)
therefor L=9+2x
48=2((2x+9)+x)
48=4x+18+2x
take 18 over
48-18=6x
30=6x
x=30/6
x=5
therefore the width is 5in.
if you doubt your answer you cn always fix it back to the eqn to test your answer